# New Types of Hydrogenlike matter Composed of Electron(s) and Meson(s)

### v1

@@ Acknowledgments
Appendix: Breit approximation, Fourier transformation, and the function of $Y(\Lambda,m,r)$
@@
$m_{i}$ and $m_{f}$ are the mass of the initial and the final states. $\mathcal{V}^{AB\to CD}(\textbf{{q}})$ and $\mathcal{M}^{AB\to CD}(\textbf{{q}})$ are the effective potential in momentum space and the scattering amplitude of $AB\to CD$ process.
The effective potentials in the momentum and coordinate space are simply connected by Fourier transformation.
@@
where $\Lambda$ is the cutoff.
When performing the Fourier transformation, we will meet the following integral
@@
In Ref. He et al. (2024); Sun et al. (2024), we has calculated the $Y(\Lambda,m,r)$ function, which can be simplified as

### v2

@@ Acknowledgments
Appendix: Breit approximation, Fourier transformation and the function of $Y(\Lambda,m,r)$
@@
$m_{i}$ and $m_{f}$ are the mass of the initial and the final states. $\mathcal{V}^{AB\to CD}(\textbf{{q}})$ and $\mathcal{M}^{AB\to CD}(\textbf{{q}})$ are the effective potential in momentum space and the scattering amplitude of $AB\to CD$ process. The effective potentials in the momentum and coordinate space are simply connected by Fourier transformation,
@@
where $\Lambda$ is the cutoff. When performing the Fourier transformation, we will meet the following integral
@@
In Ref. He:2024aej; Sun:2024wxz, we has calculated the $Y(\Lambda,m,r)$ function, which can be simplified as